40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2     Entry time: Sun Oct 21 23:11:20 2007
Author: rana 
Type: Laser 
Category: MOPA 
Subject: Laser Turn on report 
When the laser was first turned on (slowly, as per the manual), the output power was only ~14 W instead of 35.

Some tweaking of the mirrors M4 & M5 (~1 full turn each in both pitch and yaw) restored the power. The peak
power was 37.0 W as measured on the Big Head. Cool

The beam is also not centered on the monitor PD and so its more likely a small shift of some earlier
mount than such large shifts of those last two mirrors.

With the cold temperature, the power has drooped to 35.5 W. Frown
ELOG V3.1.3-